# biblatexref.bib

### v2

@@ 5. Examples
In this section, we apply Theorem 4.3 to explicit examples of homogeneous varieties. For more details on these varieties, we refer to [IMPANGA]. The projective normality hypothesis follows from [RR85]. The section-dominating collections of line bundles follow from Examples 2.5 and 2.6. We remark that we can also apply the theorem to products of these varieties.
@@ 5.1. Grassmannians
Let $A=G(k,n)=G$ be the Grassmannian of $k$-dimensional subspaces in an $n$-dimensional vector space. Consider $G$ embedded in $\mathbb{P}^{N}$ via the Plücker embedding $p:G\to\mathbb{P}^{N}$. Let $H$ be the hyperplane section in $G$ corresponding to $\mathcal{O}_{G}(1)=p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$. Let $\mathcal{E}$ be the line bundle $dH$ for some degree $d>0$. The dimension of $G$ is $\dim G=k(n-k)$. The canonical bundle of $G$ is $\omega_{G}=\mathcal{O}_{G}(-n)$.
@@
  • If $\dim G=k(n-k)\geq 4$ and $d\geq k(n-k)+n-2$, then a very general degree $d$ hypersurface $X\subset G(k,n)$ is algebraically hyperbolic.
@@ 5.3. Orthogonal Grassmannians
Let $A$ be the orthogonal Grassmannian $OG(k,n)$. Similarly, let $H$ be the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. We have $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}$ and $K_{OG}=(-n+3k-1)H$.
@@
  • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-3k-1$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k-1)}{2}+n-3k-3$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@ 5.4. Symplectic Grassmannians
Let $A=SG(k,n)$ be a symplectic Grassmannian. Let $H$ by the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. In this case, $\dim SG(k,n)=\frac{k(2n-3k+1)}{2}$ and $K_{SG}=(-n+3k-2)H$.
@@
  • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n+3k$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k+1)}{2}+n+3k-2$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@ 5.5. Flag varieties
give, via pull-back, the hyperplane classes $H_{1},\ldots,H_{m}$ that generate the Picard group of $A$. The dimension of $A$ is $D=\sum_{i=0}^{m}k_{i+1}(k_{i+1}-k_{i})$. The canonical divisor of $A$ is $K_{A}=-\sum_{i=0}^{m-1}(k_{i+2}-k_{i})H_{i+1}=-k_{2}H_{1}-(k_{3}-k_{1})H_{2}-\cdots-(n-k_{m-1})H_{m}$.
@@
  • If $\dim A=\sum_{i=0}^{m}k_{i+1}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=0}^{m}k_{j+1}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\left(\sum_{j=0}^{m}k_{j+1}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.
[…]
[title=References]

### v3

@@ 5. Examples
In this section, we apply Theorem 4.3 to explicit examples of homogeneous varieties. For more details on these varieties, we refer to [Cos18]. The projective normality hypothesis follows from [RR85]. The section-dominating collections of line bundles follow from Examples 2.5 and 2.6. We remark that we can also apply the theorem to products of these varieties.
@@ 5.1. Grassmannians
Let $A=G(k,n)$ be the Grassmannian of $k$-dimensional subspaces in an $n$-dimensional vector space. Consider $G(k,n)$ embedded in $\mathbb{P}^{N}$ via the Plücker embedding $p:G(k,n)\to\mathbb{P}^{N}$. Let $H$ be the hyperplane section in $G(k,n)$ corresponding to $\mathcal{O}_{G(k,n)}(1)=p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$. Let $\mathcal{E}$ be the line bundle $dH$ for some degree $d>0$. The dimension of the Grassmannian is $\dim G(k,n)=k(n-k)$. The tangent bundle is $T_{G(k,n)}\cong S^{*}\otimes Q$, where $S$ is the tautological bundle and $Q$ is the quotient bundle of $G(k,n)$. Then a Chern class computation gives
$c_{1}(T_{G(k,n)})=c_{1}(S^{*}\otimes Q)=nH.$
Thus, the canonical divisor is $K_{G(k,n)}=-nH$.
@@
  • If $\dim G(k,n)=k(n-k)\geq 4$ and $d\geq k(n-k)+n-2$, then a very general degree $d$ hypersurface $X\subset G(k,n)$ is algebraically hyperbolic.
@@ 5.3. Orthogonal Grassmannians
Let $A$ be the orthogonal Grassmannian $OG(k,n)$. Similarly, let $H$ be the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. We have $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}$. The orthogonal Grassmannian is the zero locus in $G(k,n)$ of a global section of $\operatorname{Sym}^{2}(S^{*})$, and we have an exact sequence of tangent bundles
$0\longrightarrow T_{OG(k,n)}\longrightarrow S^{*}\otimes Q\longrightarrow\operatorname{Sym}^{2}(S^{*})\longrightarrow 0$
So, by a Chern class computation, we have
$c_{1}(T_{OG(k,n)})=c_{1}(S^{*}\otimes Q)-c_{1}(\operatorname{Sym}^{2}(S^{*}))=(n-k-1)H.$
Thus, the canonical divisor of $OG(k,n)$ is $K_{OG}=-(n-k-1)H$.
@@
  • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-k-3$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k-1)}{2}+n-k-5$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@ 5.4. Symplectic Grassmannians
Let $A=SG(k,n)$ be a symplectic Grassmannian. Let $H$ by the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. In this case, $\dim SG(k,n)=\frac{k(2n-3k+1)}{2}$. Similarly, we have an exact sequence
$0\longrightarrow T_{SG(k,n)}\longrightarrow S^{*}\otimes Q\longrightarrow\wedge^{2}(S^{*})\longrightarrow 0,$
so we can compute
$c_{1}(T_{SG(k,n)})=c_{1}(S^{*}\otimes Q)-c_{1}(\wedge^{2}(S^{*}))=(n-k+1)H.$
Thus, the canonical divisor is $K_{SG(k,n)}=-(n-k+1)H$.
@@
  • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n-k-1$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k+1)}{2}+n-k-3$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@ 5.5. Flag varieties
give, via pull-back, the hyperplane classes $H_{1},\ldots,H_{m}$ that generate the Picard group of $A$. To fix notation, let $k_{0}=0$ and $k_{m+1}=n$. The dimension of $A$ is $\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})$. The tangent bundle of $A$ is isomorphic to $\bigoplus_{i=1}^{m}S_{k_{i}}^{*}\otimes(S_{k_{i+1}}/S_{k_{i}})$, where $S_{k_{i}}$ is the rank $k_{i}$ tautological bundle. Then, by a Chern class computation,
$c_{1}(F(k_{1},\ldots,k_{m};n))=c_{1}\left(\bigoplus_{i=1}^{m}S_{k_{i}}^{*}\otimes(S_{k_{i+1}}/S_{k_{i}})\right)=\sum_{i=1}^{m}(k_{i+1}-k_{i-1})H_{i}.$
Therefore, the canonical divisor of $F(k_{1},\ldots,k_{m};n)$ is $K_{A}=-\sum_{i=1}^{m}(k_{i+1}-k_{i-1})H_{i}=-k_{2}H_{1}-(k_{3}-k_{1})H_{2}-\cdots-(n-k_{m-1})H_{m}$.
@@
  • If $\dim A=\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.